OpenStudy (anonymous):

if f(x)=(2x+x)/(x+2), find the equation for y=f^(-1)(x)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

typo? do you really mean \[f(x)=\frac{2x+x}{x+2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but 2x+x=3x!

OpenStudy (anonymous):

why isn't it just \[f(x)=\frac{3x}{x+2}\]?

OpenStudy (anonymous):

i dont get it???